# Dance to Demise – How Massive Stars May Form Dense Circumstellar Shells Before Explosion

### v2

@@ Abstract
We investigate the evolution of red supergiant (RSG) progenitors of core-collapse (CC) supernovae (SNe) with initial masses between M⊙ focusing on the effects of enhanced mass loss due to pulsation-driven instabilities in their envelopes and subsequent dynamical ejections during advanced stages of nuclear burning. Using time-dependent mass loss from detailed MESA stellar evolution models, including a parameterized prescription for pulsation-driven superwinds and time-averaged mass loss rates attributed to resulting shock-induced ejections, we construct the circumstellar medium (CSM) before the SN explosion. We calculate resulting CSM density profiles and column densities considering the acceleration of the stellar wind. Our models produce episodes of enhanced mass loss ( M⊙ ) in the last centuries-decades before explosion forming dense CSM ( at distances ) – consistent with those inferred from multi-wavelength observations of Type II SNe such as SN 2023ixf, SN 2020ywx, SN 2017hcc, SN 2005ip and SN 1998S. The formation of such dense CS shells, within the explored range of our single star RSG models, provides a natural explanation for observed flash-ionization signatures, X-ray and radio emission, and has important implications for dust formation around Type II SNe.
@@ 1 Background
Massive stars with initial masses between M⊙ (A. Heger et al., 2003; G. Meynet & A. Maeder, 2003) are expected to end their lives in core-collapse (CC) supernovae (SNe), and are considered to be significant dust producers in galaxies (B. E. K. Sugerman et al., 2006). However, the evolution of massive stars in their late phases is poorly understood - in particular, mass loss and its impact on the final fate of these stars remains ill-constrained (N. Smith, 2014; G. Meynet et al., 2015). A number of factors have been invoked to explain mass loss from massive stars like line-driven winds, stellar rotation, radial pulsations, episodic ejections/eruptions and binarity (S. E. Woosley et al., 2002; N. Langer, 2012). All these factors result in the observed diversity of CCSNe, which are broadly classified as Type I (b/c) – which do not show any hydrogen lines in their spectra – and Type II (P/L/n/b) which do so, but with varying strengths and time evolution of light curves (A. V. Filippenko, 1997; L. Martinez et al., 2022; I. Salmaso et al., 2025).
@@
Observationally, direct identifications of SN progenitors (S. J. Smartt, 2009) provide critical information regarding the final stage of its evolution before explosion, and there exist many pre-SN images of Type II progenitors (S. J. Smartt, 2015; S. D. Van Dyk, 2017) The wealth of observations of SNe from large-scale transient surveys like the Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2019), the Asteroid Terrestrial-impact Last Alert System (ATLAS; J. L. Tonry et al., 2018) and the upcoming Legacy Survey of Space and Time (LSST; K. M. Hambleton et al., 2023), provides golden opportunities to improve our current understanding of the final stages leading to the explosive end of the lives of massive stars (C. L. Ransome & V. A. Villar, 2024; N. L. Strotjohann et al., 2024).
@@
However, several factors complicate the task of deriving reliable mass loss rates () for cool, possibly dust-enshrouded, red supergiant (RSG) progenitors of Type II SNe (J. T. van Loon et al., 2005; L. Decin, 2021). At the same time, there is a growing consensus that a high fraction of Type II SNe are preceded by phases of heavy mass loss (K.-R. Hinds et al., 2025) whose signatures can be seen from the post-explosion shock interaction with the circumstellar medium (CSM) created by the wind of the SN progenitor (P. Chandra, 2018; L. Dessart, 2024; V. V. Dwarkadas, 2025). The circumstellar densities inferred from flash-ionization spectroscopy, X-ray and radio observations of SNe like SN 1998S (N. Chugai et al., 2002), SN 2010jl (C. Fransson et al., 2014), and SN 2023ixf (K. A. Chugai et al., 2023) point toward mass loss rates of M⊙ within cm of the progenitor —- values that far exceed those from standard steady winds. The presence of narrow P Cygni absorption features, light curve plateaus with CSM interaction signatures, and late-time IR emission all corroborate the scenario of enhanced pre-SN mass loss. All these observations of post-explosion features are clearly at odds with latest mass loss prescriptions inferred from large samples of observed RSGs (E. R. Beasor et al., 2023; P. Massey et al., 2023; M. Yang et al., 2023; K. Antoniadis et al., 2024; L. Decin et al., 2024) as well as empirical mass loss rates derived from much smaller samples of stars (C. de Jager et al., 1988; H. Nieuwenhuijzen & C. de Jager, 1990) that are still used in state-of-the-art stellar evolution codes to model the RSG phase of massive stars (N. Mauron & E. Josselin, 2011; P. Massey et al., 2021).
@@
From a theoretical point of view, there have only been handful of attempts to derive mass loss prescriptions for RSGs from first principles. S.-C. Yoon & M. Cantiello (2010) proposed a pulsation-driven wind mass loss prescription for high mass RSGs with high luminosity-to-mass ratios, which they hypothesized could, in principle, achieve the high rates ( M⊙ ) required to explain observed properties of CSM around Type IIn SNe. Thereafter, Moriya, Takashi J. & Langer, Norbert (2015) followed similar principles to propose a pulsation-driven wind for very massive metal-free stars. M. Clayton (2018) further extended the pulsation-driven mass loss into the shock-dominated regime and proposed analytical mass loss rates for use in stellar evolutionary calculations. N. D. Kee et al. (2021) proposed an alternate mechanism to drive mass loss through the pressure due to characteristic turbulent velocities inferred for RSGs. Most recently, J. S. Vink & G. N. Sabhahit (2023) proposed a new mass loss prescription for RSGs as a function of the Eddington factor based on the data from M. Yang et al. (2023). However, none of these existing prescriptions by itself can fully explain all the observational constraints such as the RSG luminosity functions (P. Massey et al., 2023). Even binary interactions are estimated to overpredict luminous RSGs arising from mass accretion or mergers (E. Zapartas et al., 2025). Hence, most evolutionary codes continue to employ empirical mass loss rates (E. R. Beasor et al., 2021) which lack physical grounding and fail to reconcile with observed properties of Type II SNe.
@@ 2 Rationale of this study
The motivation behind this work is to address (a) whether a single star may incur heavy mass loss in the final stages before it explodes, that can be correlated with the presence of a dense CSM near the star, and (b) how the progenitor mass impacts the mass loss rates, CSM densities, and wind velocities. In this regard, our study accounts for progenitors of Type II supernovae, and we compare the CSM profiles of well-observed interacting SNe such as SN 2023ixf (J. E. Jencson et al., 2023; W. V. Jacobson-Galán et al., 2023; E. A. Zimmerman et al., 2024; N. Smith et al., 2023; K. A. Bostroem et al., 2024; N. A. J. et al., 2025), SN 2005ip (N. Smith et al., 2009; S. Katsuda et al., 2014), SN 1998S (N. Chugai et al., 2002; M. Pozzo et al., 2004; C. Fransson et al., 2005; J. Mauerhan & N. Smith, 2012; I. Shivvers et al., 2015), SN 2017hcc (P. Chandra et al., 2022), SN 2020ywx (R. Baer-Way et al., 2025). We argue that for single stars, enhanced pulsation near the end of their lives provides conditions favorable for escalated mass loss (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015). To demonstrate this effect, we simulate the evolution of massive stars with initial masses between and M⊙, from their main-sequence until the end of neon burning (or until the end of core O-burning, in some cases) using the MESA code, applying advanced mass loss prescriptions to incorporate effects of radial pulsations in the outer envelope of RSGs (Section 4).
The paper is arranged as follows: in Section 3, we briefly review the various mass loss mechanisms in RSG progenitors of CCSNe along with observational evidence for enhanced mass loss. In Section 4, we describe our modeling approach for incorporating enhanced mass loss in pulsating RSG progenitors followed by the model predictions for the properties of the CSM. In Section 5, we compare our model predictions with available observational estimates of mass loss rates, CSM density and column density along line of sight and discuss the implications of our model predictions for upcoming infrared observations from large surveys like the James Webb Space Telescope (JWST; J. P. Gardner et al., 2023) and the LSST in Section 6. Finally, we conclude with future directions of work in synergy with upcoming transient surveys.
@@ 3 Mass loss in RSGs
Mass loss from RSGs is a fundamental yet still uncertain ingredient in the late stages of stellar evolution (H. J. Lamers & E. M. Levesque, 2017; R. M. Humphreys, 2025; S. D. Van Dyk, 2025; J. T. van Loon, 2025; S. Ekström & C. Georgy, 2025). The physical mechanisms responsible for this process directly shape the CSM, influencing both the pre-SN environment and the post-explosion observables, including light curves, spectra, and dust formation (E. M. Levesque, 2017). Observations of Type II SNe with dense CSM (e.g., SN 2009ip, SN 2010jl, SN 2023ixf and so on) have revealed that enhanced and often eruptive mass loss can occur shortly before CC (T. Zhang et al., 2012; E. O. Ofek et al., 2013; K. A. Chugai et al., 2023). In this section, we describe potential driving mechanisms for steady line-driven winds, pulsation-driven outflows including dynamical ejections and/or eruptions, turbulent pressure and convective shocks creating extended atmosphere/chromosphere for dust-driven winds, and enhanced mass loss due to binary interaction for RSGs.
@@ 3.1 Line-driven wind
Line-driven winds are driven by the transfer of momentum from photons to metal lines in the stellar atmosphere. This mechanism is efficient in hot, compact stars e.g., OB stars (J. S. Vink, 2017), but is considerably less effective in RSGs due to their cooler photospheres ( K), where fewer atoms are ionized and UV photons are scarce (J. I. Castor et al., 1975; L. B. van den Hoek & M. A. T. Groenewegen, 1997). Typical mass loss rates from line-driven winds in RSGs range from to M⊙ (E. R. Beasor et al., 2021; J. Merritt et al., 2025), which is inadequate to explain the dense CSM observed in the majority (upto ) of all Type II SNe with early-time observations (V. Morozova et al., 2018; F. Förster et al., 2018).
@@ 3.2 Pulsation driven mass loss
Many RSGs exhibit large-amplitude radial pulsations with periods of several hundred days (M. D. Soraisam et al., 2018; Y. Ren et al., 2019), as observed in long-period variables (LPVs). Such radial motions in the H-rich envelope of RSGs are driven by the so-called -mechanism (N. Baker & R. Kippenhahn, 1965; T. Gastine & B. Dintrans, 2008) - an interplay between excess radiation pressure and gravitational pull on the outer layers of the RSG envelope. These pulsations can grow in amplitude to produce shocks at the stellar surface that levitate material above the photosphere, reducing the effective gravity and allowing other processes (e.g., radiation pressure on dust and molecules) to accelerate the material outward. However, the exact mechanism of mass loss due to the complex interplay of pulsations and convection in RSG envelopes remains ill-understood (T. Gastine & B. Dintrans, 2010), and attempts to reduce these inherently three-dimensional phenomena into one-dimensional prescriptions for use in evolutionary codes remain limited (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015; V. A. Bronner et al., 2025).
@@ 3.2.1 Dynamical Ejections
In advanced nuclear burning stages, vigorous convection and shell flashes can lead to hydrodynamical instabilities and formation of shock waves in the outer envelopes of RSGs. If these internal shocks propagate outward and reach the stellar surface, they can trigger impulsive ejections of envelope material. These “dynamical ejections” are thought to be responsible for the formation of dense, compact shells detected in SN progenitors such as SN 2009ip (J. C. Mauerhan et al., 2013; N. Smith, 2014). Simulations suggest that mass loss rates in such events can temporarily exceed M⊙ , lasting sometimes for days to weeks or even sustain repeatedly for centuries (M. Clayton, 2018).
@@ 3.2.2 Eruptive mass loss
Eruptive mass loss may occur through surface instabilities driven by opacity changes, pulsationally enhanced convective transport, wave-driven energy deposition or super-Eddington luminosities in RSG envelopes (E. Quataert & J. Shiode, 2012; J. Fuller, 2017; S. J. Cheng et al., 2024). These eruptions resemble luminous blue variable (LBV) behavior, though they are seen in cooler RSGs as well. The pre-SN outbursts seen in light echoes (e.g., SN 2013cu) are likely signatures of such events, with inferred – M⊙ over timescales of months to years (G. Gräfener & J. S. Vink, 2016). Furthermore, statistical analysis of luminous stars in the Milky Way (C. S. Kochanek, 2011) and the study of Carinae analogs in nearby galaxies (R. Khan et al., 2015) reveal the presence of massive shells which hint towards instabilities triggered at the onset of C-burning leading to eruptions (R. Margutti et al., 2017).
@@ 3.3 Turbulent pressure and Extended Chromosphere
N. D. Kee et al. (2021) have recently presented a model showing how turbulent pressure inside the RSG envelope is sufficient to expand the atmosphere to an extent that a wind can be launched. Their analytic mass loss prescription for observationally inferred turbulent velocities (B. Arroyo-Torres et al., 2015) predicts that stars above M⊙ leave the RSG phase to evolve into blue supergiants and hence would not explode as TypeII P/L SNe – in agreement with observations (S. J. Smartt et al., 2009) and in contrast to results obtained using empirical mass loss rates (C. de Jager et al., 1988; E. R. Beasor et al., 2021) in which massive stars below M⊙end their lives as RSGs.
Along similar lines, J. Fuller & D. Tsuna (2024) proposed the existence of a dense chromosphere supported by outgoing shocks near the photosphere created from strong convective motions inside the RSG envelopes. The chromosphere in their model extends to several stellar radii111In J. Fuller & D. Tsuna (2024), the stellar radius is the outer boundary of the relaxed stellar model, located at a very low optical depth so that their models give a smooth and realistic transition to the circumstellar environment., out to distances where dust formation can drive the wind launched by the radiation pressure of the star (E. Sedlmayr & C. Dominik, 1995). The mass loss rates predicted from their model are similar to recent observations, with increasing values for more luminous RSGs, and can potentially explain the early rise in light curves of Type IIP SNe without invoking pre-explosion enhanced mass loss.

### v3

@@ Abstract
We investigate the evolution of red supergiant (RSG) progenitors of core-collapse supernovae (SNe) with initial masses between and M⊙, focusing on the effects of enhanced mass loss due to pulsation-driven instabilities in their envelopes and subsequent dynamical ejections during advanced stages of nuclear burning. Using time-dependent mass loss from detailed Modules for Experiments in Stellar Astrophysics (MESA) stellar evolution models, including a parameterized prescription for pulsation-driven superwinds and time-averaged mass loss rates attributed to resulting shock-induced ejections, we construct the circumstellar medium (CSM) before the SN explosion. We calculate resulting CSM density profiles and column densities considering the acceleration of the stellar wind. Our models produce episodes of enhanced mass loss ( M⊙ ) in the last centuries-decades before explosion forming dense CSM ( at distances ) – consistent with those inferred from multi-wavelength observations of Type II SNe such as SN 2023ixf, SN 2020ywx, SN 2017hcc, SN 2005ip and SN 1998S. The formation of such dense CS shells, within the explored range of our single star RSG models, provides a natural explanation for observed flash-ionization signatures, X-ray and radio emission, and has important implications for dust formation around Type II SNe.
@@ 1 Background
Massive stars with initial masses between M⊙ (Heger_2003; 2003A&A...404..975M) are expected to end their lives in core-collapse (CC) supernovae (SNe), and are considered to be significant dust producers in galaxies (sug06). However, the evolution of massive stars in their late phases is poorly understood - in particular, mass loss and its impact on the final fate of these stars remains ill-constrained (2014ARA&A..52..487S; 2015A&A...575A..60M). A number of factors have been invoked to explain mass loss from massive stars like line-driven winds, stellar rotation, radial pulsations, episodic ejections/eruptions and binarity (2002RvMP...74.1015W; 2012ARA&A..50..107L). All these factors result in the observed diversity of CCSNe, which are broadly classified as Type I (b/c) – which do not show any hydrogen lines in their spectra – and Type II (P/L/n/b) which do so, but with varying strengths and time evolution of light curves (1997ARA&A..35..309F; 2022A&A...660A..40M; 2025A&A...695A..29S).
@@
Observationally, direct identifications of SN progenitors (smartt_2009) provide critical information regarding the final stage of its evolution before explosion, and there exist many pre-SN images of Type II progenitors (2015PASA...32...16S; 2017hsn..book..693V) The wealth of observations of SNe from large-scale transient surveys like the Zwicky Transient Facility (ZTF; 2019PASP..131a8002B), the Asteroid Terrestrial-impact Last Alert System (ATLAS; 2018PASP..130f4505T) and the upcoming Legacy Survey of Space and Time (LSST; 2023PASP..135j5002H), provides golden opportunities to improve our current understanding of the final stages leading to the explosive end of the lives of massive stars (2024arXiv240910596R; 2024ApJ...960...72S).
@@
However, several factors complicate the task of deriving reliable mass loss rates () for cool, possibly dust-enshrouded, red supergiant (RSG) progenitors of Type II SNe (2005A&A...438..273V; 2021ARA&A..59..337D). At the same time, there is a growing consensus that a high fraction of Type II SNe are preceded by phases of heavy mass loss (10.1093/mnras/staf888) whose signatures can be seen from the post-explosion shock interaction with the circumstellar medium (CSM) created by the wind of the SN progenitor (2018SSRv..214...27C; 2024arXiv240504259D; universe11050161). The circumstellar densities inferred from flash-ionization spectroscopy, X-ray and radio observations of SNe like SN 1998S (10.1046/j.1365-8711.2002.05086.x), SN 2010jl (2014ApJ...797..118F), and SN 2023ixf (2023ApJ...956L...5B) point toward mass loss rates of M⊙ within cm of the progenitor —- values that far exceed those from standard steady winds. The presence of narrow P Cygni absorption features, light curve plateaus with CSM interaction signatures, and late-time IR emission all corroborate the scenario of enhanced pre-SN mass loss. All these observations of post-explosion features are clearly at odds with latest mass loss prescriptions inferred from large samples of observed RSGs (10.1093/mnras/stad1818; 2023ApJ...942...69M; 2023A&A...676A..84Y; antoniadis_2024; 2024A&A...681A..17D) as well as empirical mass loss rates derived from much smaller samples of stars (1988A&AS...72..259D; 1990A&A...231..134N) that are still used in state-of-the-art stellar evolution codes to model the RSG phase of massive stars (2011A&A...526A.156M; 2021ApJ...922..177M).
@@
From a theoretical point of view, there have only been handful of attempts to derive mass loss prescriptions for RSGs from first principles. 2010ApJ...717L..62Y proposed a pulsation-driven wind mass loss prescription for high mass RSGs with high luminosity-to-mass ratios, which they hypothesized could, in principle, achieve the high rates ( M⊙ ) required to explain observed properties of CSM around Type IIn SNe. Thereafter, ml2015 followed similar principles to propose a pulsation-driven wind for very massive metal-free stars. clayton2018a further extended the pulsation-driven mass loss into the shock-dominated regime and proposed analytical mass loss rates for use in stellar evolutionary calculations. 2021A&A...646A.180K proposed an alternate mechanism to drive mass loss through the pressure due to characteristic turbulent velocities inferred for RSGs. Most recently, 2023A&A...678L...3V proposed a new mass loss prescription for RSGs as a function of the Eddington factor based on the data from 2023A&A...676A..84Y. However, none of these existing prescriptions by itself can fully explain all the observational constraints such as the RSG luminosity functions (2023ApJ...942...69M). Even binary interactions are estimated to overpredict luminous RSGs arising from mass accretion or mergers (2025A&A...697A.167Z). Hence, most evolutionary codes continue to employ empirical mass loss rates (2021ApJ...922...55B) which lack physical grounding and fail to reconcile with observed properties of Type II SNe.
@@ 2 Rationale of this study
The motivation behind this work is to address (a) whether a single star may incur heavy mass loss in the final stages before it explodes, that can be correlated with the presence of a dense CSM near the star, and (b) how the progenitor mass impacts the mass loss rates, CSM densities, and wind velocities. In this regard, our study accounts for progenitors of Type II supernovae, and we compare the CSM profiles of well-observed interacting SNe such as SN 2023ixf (2023ApJ...952L..30J; 2023ApJ...954L..42J; 2024Natur.627..759Z; 2023ApJ...956...46S; 2024ApJ...973L..47B; 2025ApJ...985...51A), SN 2005ip (Smith_2009; Katsuda_2014), SN 1998S (10.1046/j.1365-8711.2002.05086.x; 2004MNRAS.352..457P; Fransson_2005; 2012MNRAS.424.2659M; 2015ApJ...806..213S), SN 2017hcc (2022MNRAS.517.4151C), SN 2020ywx (Baer-Way_2025). We argue that for single stars, enhanced pulsation near the end of their lives provides conditions favorable for escalated mass loss (2010ApJ...717L..62Y; ml2015). To demonstrate this effect, we simulate the evolution of massive stars with initial masses between and M⊙, from their main-sequence until the end of neon burning (or until the end of core O-burning, in some cases) using the MESA code, applying advanced mass loss prescriptions to incorporate effects of radial pulsations in the outer envelope of RSGs (Section 4).
The paper is arranged as follows: in Section 3, we briefly review the various mass loss mechanisms in RSG progenitors of CCSNe along with observational evidence for enhanced mass loss. In Section 4, we describe our modeling approach for incorporating enhanced mass loss in pulsating RSG progenitors followed by the model predictions for the properties of the CSM. In Section 5, we compare our model predictions with available observational estimates of mass loss rates, CSM density and column density along line of sight and discuss the implications of our model predictions for upcoming infrared observations from large surveys like the James Webb Space Telescope (JWST; Gardner_2023) and the LSST in Section 6. Finally, we conclude with future directions of work in synergy with upcoming transient surveys.
@@ 3 Mass loss in RSGs
Mass loss from RSGs is a fundamental yet still uncertain ingredient in the late stages of stellar evolution (10.1088/978-0-7503-1278-3ch15; galaxies13020025; galaxies13020033; galaxies13040072; galaxies13040081). The physical mechanisms responsible for this process directly shape the CSM, influencing both the pre-SN environment and the post-explosion observables, including light curves, spectra, and dust formation (10.1088/978-0-7503-1329-2ch4). Observations of Type II SNe with dense CSM (e.g., SN 2009ip, SN 2010jl, SN 2023ixf and so on) have revealed that enhanced and often eruptive mass loss can occur shortly before CC (2012AJ....144..131Z; 2013ApJ...768...47O; 2023ApJ...956L...5B). In this section, we describe potential driving mechanisms for steady line-driven winds, pulsation-driven outflows including dynamical ejections and/or eruptions, turbulent pressure and convective shocks creating extended atmosphere/chromosphere for dust-driven winds, and enhanced mass loss due to binary interaction for RSGs.
@@ 3.1 Line-driven wind
Line-driven winds are driven by the transfer of momentum from photons to metal lines in the stellar atmosphere. This mechanism is efficient in hot, compact stars e.g., OB stars (2017RSPTA.37560269V), but is considerably less effective in RSGs due to their cooler photospheres ( K), where fewer atoms are ionized and UV photons are scarce (1975ApJ...195..157C; 1997A&AS..123..305V). Typical mass loss rates from line-driven winds in RSGs range from to M⊙ (2021ApJ...922...55B; 2025arXiv250717052M), which is inadequate to explain the dense CSM observed in the majority (upto ) of all Type II SNe with early-time observations (2018ApJ...858...15M; 2018NatAst2.10.808).
@@ 3.2 Pulsation driven mass loss
Many RSGs exhibit large-amplitude radial pulsations with periods of several hundred days (2018ApJ...859...73S; 2019ApJS..241...35R), as observed in long-period variables (LPVs). Such radial motions in the H-rich envelope of RSGs are driven by the so-called -mechanism (1965ApJ...142..868B; 2008A&A...484...29G) - an interplay between excess radiation pressure and gravitational pull on the outer layers of the RSG envelope. These pulsations can grow in amplitude to produce shocks at the stellar surface that levitate material above the photosphere, reducing the effective gravity and allowing other processes (e.g., radiation pressure on dust and molecules) to accelerate the material outward. However, the exact mechanism of mass loss due to the complex interplay of pulsations and convection in RSG envelopes remains ill-understood (2010Ap&SS.328..245G), and attempts to reduce these inherently three-dimensional phenomena into one-dimensional prescriptions for use in evolutionary codes remain limited (2010ApJ...717L..62Y; ml2015; 2025A&A...703A..61B; 2025MNRAS.543.3929S).
@@ 3.2.1 Dynamical Ejections
In advanced nuclear burning stages, vigorous convection and shell flashes can lead to hydrodynamical instabilities and formation of shock waves in the outer envelopes of RSGs. If these internal shocks propagate outward and reach the stellar surface, they can trigger impulsive ejections of envelope material. These “dynamical ejections” are thought to be responsible for the formation of dense, compact shells detected in SN progenitors such as SN 2009ip (2013MNRAS.430.1801M; 2014ARA&A..52..487S). Simulations suggest that mass loss rates in such events can temporarily exceed M⊙ , lasting sometimes for days to weeks or even sustain repeatedly for centuries (clayton2018a).
@@ 3.2.2 Eruptive mass loss
Eruptive mass loss may occur through surface instabilities driven by opacity changes, pulsationally enhanced convective transport, wave-driven energy deposition or super-Eddington luminosities in RSG envelopes (2012MNRAS.423L..92Q; 2017MNRAS.470.1642F; Cheng_2024). These eruptions resemble luminous blue variable (LBV) behavior, though they are seen in cooler RSGs as well. The pre-SN outbursts seen in light echoes (e.g., SN 2013cu) are likely signatures of such events, with inferred – M⊙ over timescales of months to years (2016MNRAS.455..112G). Furthermore, statistical analysis of luminous stars in the Milky Way (2011ApJ...743...73K) and the study of Carinae analogs in nearby galaxies (2015ApJ...815L..18K) reveal the presence of massive shells which hint towards instabilities triggered at the onset of C-burning leading to eruptions (Margutti_2017).
@@ 3.3 Turbulent pressure and Extended Chromosphere
2021A&A...646A.180K have recently presented a model showing how turbulent pressure inside the RSG envelope is sufficient to expand the atmosphere to an extent that a wind can be launched. Their analytic mass loss prescription for observationally inferred turbulent velocities (2015A&A...575A..50A) predicts that stars above M⊙ leave the RSG phase to evolve into blue supergiants and hence would not explode as TypeII P/L SNe – in agreement with observations (2009MNRAS.395.1409S) and in contrast to results obtained using empirical mass loss rates (1988A&AS...72..259D; 2021ApJ...922...55B) in which massive stars below M⊙end their lives as RSGs.
Along similar lines, 2024OJAp....7E..47F proposed the existence of a dense chromosphere supported by outgoing shocks near the photosphere created from strong convective motions inside the RSG envelopes. The chromosphere in their model extends to several stellar radii 111In 2024OJAp....7E..47F, the stellar radius is the outer boundary of the relaxed stellar model, located at a very low optical depth so that their models give a smooth and realistic transition to the circumstellar environment., out to distances where dust formation can drive the wind launched by the radiation pressure of the star (1995SSRv...73..211S). The mass loss rates predicted from their model are similar to recent observations, with increasing values for more luminous RSGs, and can potentially explain the early rise in light curves of Type IIP SNe without invoking pre-explosion enhanced mass loss.